OpenStudy (anonymous):

N ..... F(n) 1 4 2 11 3 18 4 25 5 32 Which function represents this data

OpenStudy (anonymous):

Do you have choices? Like A, B, C, D?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Can you give them?

OpenStudy (anonymous):

a. f (n)= 3n+1 b. f (n)= 5n+3 c. f(n)= 6n-1 d. f(n)= 7n-3

OpenStudy (anonymous):

It's D, just plug the first number into N and the second number should come out as f(n) if it is the right answer.